2Na+Cl2→2NaCl

When 10.0 g of Cl2 is used, the reaction produced 12.5 grams NaCl. What is the % yield of the reaction?


Use the following molar masses: Cl2=70g/mol and NaCl=58g/mol.



1
Expert's answer
2021-01-22T06:22:14-0500

2Na + Cl2 → 2NaCl

The theoretical yield of the reaction:

m(NaCl) = m(Cl2) × 2 Mr(NaCl) / Mr(Cl2) = 10.0 g × 2 × 58 g/mol / 70 g/mol = 16.6 g

As theoretical yield is 16.6 g and actual yield is 12.5 g NaCl, the percent yield equals:

w% = (actual yield / theoretical yield) × 100% = 12.5 g / 16.6 g = 75.3%


Answer: 75.3%